# Boosting Safe Human-Robot Collaboration Through Adaptive Collision Sensitivity

### v1

@@ Abstract
What is considered safe for a robot operator during physical human-robot collaboration (HRC) is specified in corresponding HRC standards (e.g., the European ISO/TS 15066). The regime that allows collisions between the moving robot and the operator, called Power and Force Limiting (PFL), restricts the permissible contact forces. Using the same fixed contact thresholds on the entire robot surface results in significant and unnecessary productivity losses, as the robot needs to stop even when impact forces are within limits. Here we present a framework for setting the protective skin thresholds individually for different parts of the robot body and dynamically on the fly, based on the effective mass of each robot link and the link velocity. We perform experiments on a 6-axis collaborative robot arm (UR10e) completely covered with a sensitive skin (AIRSKIN) consisting of eleven individual pads. On a mock pick-and-place scenario with both transient and quasi-static collisions, we demonstrate how skin sensitivity influences the task performance and exerted force. We show an increase in productivity of almost from the most conservative setting of collision thresholds to the most adaptive setting, while ensuring safety for human operators. The method is applicable to any robot for which the effective mass can be calculated.
@@ I Introduction
Robots are leaving safety fences and start to share workspace or even living space with humans. As they leave controlled environments and enter domains that are much less structured, they need to dynamically adapt to unpredictable interactions and guarantee safety at every moment. There has been rapid development in this field of physical human-robot interaction (pHRI) or human-robot collaboration (HRC) in the last decade (see [1, 2, 3] for surveys), a growing market of collaborative robots, along with the introduction and revision of new safety standards ([4, 5, 6]). According to [6], there are two main ways of satisfying the safety requirements when a human physically collaborates with a robot: (i) Power and Force Limiting (PFL) and (ii) Speed and Separation Monitoring (SSM). For PFL, physical contacts with a moving robot are allowed, but the forces, pressures, and energy absorbed during a collision need to be within part-specific limits of the human body. This translates into lightweight structure, soft padding,
@@
and no pinch points on the robot side, in combination with robot interaction control methods for the post-impact phase: collision detection and response relying on motor load measurements, force/torque, or joint torque sensing (e.g., [7]). However, the speed and payload of robots that comply with this safety requirement is limited.
@@
Furthermore, in practice, exploiting the full potential of the PFL regime, i.e., stopping the robot only if the prescribed force threshold is exceeded, is difficult or impossible. According to [8], a collision needs to be 1) detected, 2) isolated (which robot link and which location on the link), 3) identified (estimating the external torque or force), 4) classified, and 5) then a robot reaction is triggered – see the bottom part of Fig. 2. The accuracy or even the possibility of performing some of these steps depends on the type of contact (single or multiple), the contact location, and in particular on the sensory equipment of the robot. Electronic skins covering large parts of the robot surface [9, 10, 11, 12] facilitate detection and isolation. For identification, force/torque or joint torque sensors together with accurate models of the robot dynamics are the best. However, correct estimation of forces, pressures, and energies on the body of the human operator is not possible for several reasons. First, only the first very brief evolution of the collision is used for collision identification. According to PFL, the first should be considered for transient contact and even longer for quasi-static contact. Second, even if external force on the robot body can be estimated, applying Newton’s third law and assuming that the same force acts on the human body is problematic. Third, without additional visual sensors, the robot does not know with which part of the body it is colliding (different limits apply and different tissue stiffness is needed to model the collision). In practice, collaborative robot applications are thus configured too conservatively, i.e. stopping the robot as soon as a collision is detected. Frequent stops of the robot then negatively impact productivity.
@@
In this work, we present a new method that aims to maximize productivity in the PFL regime as illustrated in Fig. 2. The main idea is to perform only collision detection and isolation, which are relatively easy and can be performed fast, and trigger a robot reaction—if necessary. The collision is detected and isolated using a pressure-sensitive electronic skin (AIRSKIN) covering the entire robot body. The key innovation here is that we online estimate the collision force from the robot velocity and effective mass111Based on the simple formula from [6] extended by the more accurate effective mass calculation involving the position in the workspace and the collision direction [13]. and then use that estimate to set a threshold to compare with the pressure value measured by the protective skin. If the estimated force is low, the robot will react by continuing the execution of the task.
@@
Our recent work on this topic [15] introduced this idea, but was largely based on simulation and featured only preliminary real robot experiments. Here, we present systematic experiments on a real robot by utilizing a robust environment allowing for repeatable measurements. In addition, the skin thresholds have been redesigned (5 thresholds compared to 3 in [15] with a better selection of these thresholds) and more repetitions. The new experimental evaluation newly contains not only transient but also quasi-static (clamping) contacts.
@@ II Method
The core of this work is the evaluation of real-world impacts in a setup that simulated human-robot cooperation with several obstacles during a task performed by the robot. This section describes the theoretical background, setup, robot, and skin used during the experiments.
II-A Robot, Skin and Software
The robot used in this work is a 6- Degree of Freedom (DoF) robotic manipulator UR10e. The arm itself is designed for collaborative purposes. Our robot is further equipped with robotic skin AIRSKIN222https://www.airskin.io/. The skin is essential for our experiments. We are able to read the pressure values (in Pascals) from individual skin parts (also called pads; 11 of them placed on our robot) at approximately .
Both the robot and the skin are integrated and interconnected using Robot Operating System (ROS) middleware. During the task performed by the robot (see next Section), we utilize control in joint velocity space to allow for quick reactions to impacts.
II-B Task and Environment
@@
The environment (Fig. 3) we developed to perform the experiments consists of three buckets (B1, B2, B3) hung on a rope from the ceiling. The buckets are filled with small stones with a total weight of . The weight was selected according to the average weight of the whole human arm from ISO/TS 15066[6]. The buckets can freely move in the space, thus simulating the impact of the robot with the human arm. The impacts with buckets are transient, that is, dynamic impacts in which a human body part is impacted by a moving robot and the human can recoil or retract [6]. The setup also includes CBSF-75-Basic certified measuring device for validation of collaborative tasks with the range of . This device is firmly attached to the table and is used to simulate a clamping scenario. This impact is quasi-static, that is, impact of a human body part with the robot, when the robot clamps or crushes the human for a prolonged period of time [6].
@@
To be able to evaluate the effect of different collision threshold settings on performance, we define a mock pick-and-place task – see the accompanying video. The direction of the movements is depicted in Fig. 3. The order of movements starts from the top green line (movement in the y-axis of the robot) and collision with the first bucket B1, continues with slight downward movement (negative z-axis) and movement in the robot’s x-axis. Then it continues with impact to B2. The robot then goes again downward to collide with the measuring device, and returns up after the collision. Finally, the robot moves in its negative y-axis and collides with bucket B3.
III Skin Sensitivity Thresholds
The initial step in the pipeline is to select the individual sensitivity thresholds of our robotic skin. We decided to select five different sensitivity thresholds, numbered from 0 (the most sensitive one) to 4 (the least sensitive). The thresholds are represented as a percentage value of the pressure range in our application and were set individually for individual pads on the robot. We measured the pressure range for each pad by hitting obstacles. Then, we selected the most sensitive threshold to be robust (a low number of false negative detections), while being sensitive enough to perceive even gentle contact. The other thresholds are set as 25, 75, 95, and of the pressure range. The reason we also selected is to count for the cases when a collision is outside the distribution of the ranges we measured (which is in theory possible, as we measured each range only five times).
IV Threshold Setting Strategies
The core components are the strategies to set sensitivity thresholds for individual skin parts. We can divide our strategies to two main section: (i) uniform (static) distribution of sensitivity thresholds; (ii) dynamic distribution.
The first option is straightforward. We set all the pads on the robot to the same (uniform) sensitivity thresholds. As we have five different thresholds, we tested five different static settings based on the threshold—Uniform 0, Uniform 1, Uniform 2, Uniform 3, Uniform 4.
The dynamic setting is more complicated and is divided into two methods.
IV-1 Dynamic – Half of Moving Mass (HMM)
@@
The first method is based purely on equation from ISO/TS 15066 [6], that defines the maximal allowed velocity given permissible force limit as
@@
where is the spring constant of impacted human body, is the effective mass of the robot, is the effective mass of the impacted human body, is the Cartesian velocity of the robot link and is the maximal permissible force. The force can be computed as
@@
The spring constant is fixed in our experiments to corresponding to the spring constant of the back of the non-dominant hand as suggested by ISO/TS 15066. We selected the mass as (and we made our buckets weigh the same) based on the average weight of human arm defined also in [6]. The mass is in ISO/TS 15066 defined as Half of Moving Mass (HMM) of the robot. In our experiments, this weight is computed half of the mass of the sum of weights of links up to the link with skin part in collision. This mass is the main difference from the second dynamic policy.
IV-2 Dynamic – Effective Mass (EM)
@@
This policy also uses Eq. 2 to compute the permissible force. However, now is actually the Effective Mass (EF) of the robot. EF is a property of the robot (each point of the robot), which can be computed using the dynamics of the robot [13]. The EF is always computed in a given direction. We consider this direction to be the direction of the link on which is the skin pad that detected collision. Using this direction , one can compute the effective mass as
@@
This property of robots is well known in the community, but is arguably still not used often. The reason is its sensitivity to proper dynamic model of the robot and instability in some configurations (mainly near the base of the robot; see [16] for discussion on the topic). We provide some examples in Fig. 4, where the EF in every configuration of our task is shown for three skin pads (the pads are in collision with any of the obstacles). We can see that the EF is in most cases lower than HMM, which was our motivation to use this property in our dynamic threshold setting policy. However, we can also see that it is higher in the beginning of the task for one of the pads. The given skin part (green pad in Fig. 1) is in the begging of the movement very close to the robot base and is moving in such a way that it realistically cannot exert any force to obstacles (see the attached video). Although it is important to keep this in mind, we still believe that the utilization of EF could have a high impact not only on Human-Robot Interaction (HRI) research.
Both of the dynamic methods have the same concept of setting of thresholds based on the estimated force. We set the thresholds from force by discretizing the possible force into five intervals. The intervals are different for transient and quasi-static contacts, as based on [6] the limits for permissible force are not the same— and for quasi-static and transient contacts, respectively. Thus based on these limits, we set the five threshold from forces as
@@
V Experiments and Results

### v2

# Adaptive Collision Sensitivity for Efficient and Safe Human-Robot Collaboration

@@ Abstract
What is considered safe for a robot operator during physical human-robot collaboration (HRC) is specified in corresponding HRC standards (e.g., ISO/TS 15066). The regime that allows collisions between the moving robot and the operator, called Power and Force Limiting (PFL), restricts the permissible contact forces. Using the same fixed contact thresholds on the entire robot surface results in significant and unnecessary productivity losses, as the robot needs to stop even when impact forces are within limits. Here we present a framework that decides whether the robot should interrupt or continue its motion based on estimated collision force computed individually for different parts of the robot body and dynamically on the fly, based on the effective mass of each robot link and the link velocity. We performed experiments on simulated and real 6-axis collaborative robot arm (UR10e) with sensitive skin (AIRSKIN) for collision detection and isolation. To demonstrate the generality of our method, we added experiments on the simulated KUKA LBR iiwa robot, where collision detection and isolation draws on joint torque sensing. On a mock pick-and-place scenario with both transient and quasi-static collisions, we demonstrate how sensitivity to collisions influences the task performance and number of stops. We show an increase in productivity over 45% from using the standard approach that interrupts the tasks during every collision. While reducing the cycle time and the number of interruptions, our framework also ensures the safety of human operators. The method is applicable to any robot for which the effective mass can be calculated.
@@ I Introduction
Robots are leaving safety fences and start to share workspace or even living space with humans. As they leave controlled environments and enter domains that are much less structured, they need to dynamically adapt to unpredictable interactions and guarantee safety at every moment. There has been rapid development in this field of physical human-robot interaction (pHRI) or human-robot collaboration (HRC) in the last decade (see [1, 2, 3] for surveys), a growing market of collaborative robots, along with the introduction and revision of new safety standards ([4, 5, 6]). According to [6], there are two main ways of satisfying the safety requirements when a human physically collaborates with a robot: (i) Power and Force Limiting (PFL)
@@
and (ii) Speed and Separation Monitoring (SSM). For PFL, physical contacts with a moving robot are allowed, but the forces, pressures, and energy absorbed during a collision need to be within part-specific limits of the human body. This translates into lightweight structure, soft padding, and no pinch points on the robot side, in combination with robot interaction control methods for the post-impact phase: collision detection and response relying on motor load measurements, force/torque, or joint torque sensing (e.g., [8]). However, the speed and payload of robots that comply with this safety requirement is limited.
@@
Furthermore, in practice, exploiting the full potential of the PFL regime, i.e., stopping the robot only if the prescribed force threshold is exceeded, is difficult or impossible. According to [7], a collision needs to be 1) detected, 2) isolated (which robot link)), 3) identified (estimating the external torque or force), 4) classified, and 5) then a robot reaction is triggered—see Fig. 2. The accuracy or even the possibility of performing some of these steps depends on the type of contact (single or multiple), the contact location, and in particular on the sensory equipment of the robot. Electronic skins covering large parts of the robot surface [9, 10, 11, 12] facilitate detection and isolation. For identification, force/torque or joint torque sensors together with accurate models of the robot dynamics are the best. However, correct estimation of forces, pressures, and energies on the body of the human operator is not possible for several reasons. First, only the first very brief evolution of the collision is used for collision identification. According to PFL, the first should be considered for transient contact and even longer for quasi-static contact. Second, even if external force on the robot body can be estimated, applying Newton’s third law and assuming that the same force acts on the human body is problematic. Third, without additional visual sensors, the robot does not know with which part of the body it is colliding (different limits apply and different tissue stiffness is needed to model the collision). In practice, collaborative robot applications are thus configured too conservatively, i.e., stopping the robot as soon as a collision is detected. Frequent stops of the robot then negatively impact productivity.
@@
In this work, we present a new method that aims to maximize productivity in the PFL regime as illustrated in Fig. 2. The idea is to perform collision detection and isolation using a robotic skin (or joint-torque sensors), “identify” the force using a simple equation and trigger a robot reaction. We tested the method on two robots, and the collision is detected and isolated using a pressure-sensitive electronic skin (AIRSKIN) covering the entire robot body of the UR10e robot (both real and simulated), or using joint-torque sensors of the KUKA LBR iiwa 7 robotic arm (only simulated). The key innovation here is that we online estimate the collision force from the robot velocity and effective mass and then use that estimate to decide how to react to the collision. If the estimated force is low, the robot will react by continuing the execution of the task. If the estimated impact force is high, the robot will stop.
@@
Our recent work on this topic [14] introduced this idea, but was largely based on simulation, featured only preliminary real robot experiments and was not safety focused. Here, we present systematic experiments on a real robot by utilizing a robust environment allowing for repeatable measurements. In addition, the collision handling has been redesigned to be safer. Previously, the collision was detected only if the pressure applied to the robot was higher than a predefined threshold. Now, every collision is detected, and only the reaction is based on the estimated force. The new experimental evaluation contains not only transient but also quasi-static (clamping) contacts. We added a second robot with different dimensions and physical properties, and different collision-sensing technology. Finally, all experiments were performed with multiple Cartesian velocities.
@@ II Method
The core of this work is the evaluation of simulated and real-world impacts in a setup that simulated human-robot cooperation with several obstacles during a task performed by the robot. This section describes the theoretical background, setup, and robots used in the experiments.
II-A Task and Environment
@@
The environment (Fig. 3) we prepared to carry out the experiments consists of three buckets (B1, B2, B3) hung on a rope from the ceiling. The buckets are filled with small stones with a total weight of . The weight was selected according to the average weight of the whole human arm from ISO/TS 15066[6]. The buckets can freely move in space, thus simulating the impact of the robot with the human arm. The impacts with buckets are transient, that is, dynamic impacts in which a human body part is impacted by a moving robot and the human can recoil or retract [6]. The setup also includes the CBSF-75-Basic certified measuring device for validation of collaborative tasks with the range of . This device is firmly attached to the table and is used to simulate a clamping scenario. This impact is quasi-static, that is, impact of a human body part with the robot, when the robot clamps or crushes the human for a prolonged period of time [6].
@@
To be able to evaluate the effect of different collision threshold settings on performance, we define a mock pick-and-place task—see the video at https://youtu.be/UyOa--6Jpj4. The direction of the movements is depicted in Fig. 3. The order of movements starts from the top green line (movement in the y-axis of the robot) and collision with the first bucket B1, continues with slight downward movement (negative z-axis) and movement in the robot’s x-axis. Then it continues with impact to B2. The robot then goes again downward to collide with the measuring device, and returns up after the collision. Finally, the robot moves in its negative y-axis and collides with bucket B3.
II-B Robot, Skin and Software
We have used two robots.
Universal Robots UR10e manipulator with Airskin. This robot is a 6- Degree of Freedom (DoF) robotic manipulator and the arm itself is designed for collaborative purposes. Our robot is further equipped with robotic skin AIRSKIN111https://www.airskin.io/. Using a special Airskin maintenance controller, the pressure values from individual skin parts (also called pads; 11 of them placed on our robot) can be retrieved at approximately . Different from standard Airskin deployment in the industry where any detected collision stops the robot, here we use the pressure sensor for collision detection and isolation (which robot link collided) and choose to stop or continue using our algorithm.
A real and simulated version of the robot was used.
KUKA LBR iiwa 7. The second robot used is also collaborative. It is equipped with joint torque sensors in each of its joints and can also estimate the external torques, allowing collision detection.
A simulated version of the robot was used.
Both robots are integrated and controlled using the Robot Operating System (ROS) middleware. The task performed by the robots (see Section II-A) is defined in the Cartesian space, but we utilize control in the joint velocity space as it gives us more control over the movements and contact reactions.
Simulation. We used simulators that we developed (available at https://rustlluk.github.io/adaptive-sensitivity/). For both robots, the underlying simulation of physics is achieved with PyBullet [15]. The simulator also contains the buckets and force-measuring device with realistic physical behavior allowing us to do a higher number of experiments than in the real world.
II-C Collision Handling Strategies
We employed three different strategies to handle collision events: (i) FACTORY; (ii) FIXED MASS; (iii) ADAPTIVE MASS.
II-C1 FACTORY
The first option is common (“industrial”) strategy used with collaborative robots in industry: Whenever a contact is detected, the robot is stopped immediately. This serves as a baseline to gauge how much productivity can be gained when a better collision reaction strategy is used.
Dynamic collision handling strategies. The two strategies described below (FIXED MASS and ADAPTIVE MASS) are dynamic, that is, the robot reacts to a collision by stopping or continuing based on properties computed on the fly.
@@ II-C1 FACTORY
Both methods are based on equation from ISO/TS 15066 [6], which defines the maximal allowed robot velocity given the permissible collision force limit as
@@
where is the spring constant of impacted human body, is the Effective Mass (EF) of the robot, is the EF of the impacted human body, is the Cartesian velocity of the robot link and is the maximal permissible force.
For this work, we need to rearrange the equation and express
We evaluate an instance of this equation for every robot link on the run, plugging in the instantaneous velocity of every robot link. If collision of a particular robot link is detected, the equation is used to estimate the impact force.
@@
The spring constant is fixed in our experiments to corresponding to the spring constant of the back of the non-dominant hand as suggested by ISO/TS 15066. We selected the mass as (and we made the hanging buckets weigh the same) based on the average weight of human arm defined also in [6].
The two methods then differ in how they determine the robot mass .
II-C2 FIXED MASS
The mass is in ISO/TS 15066 defined as Half of Moving Mass (HMM) of the robot. In our experiments, this weight is computed as half of the sum of masses of links up to the link in collision.
II-C3 ADAPTIVE MASS
@@
This policy also uses Eq. 2 to compute the permissible force. However, now is actually the real Effective Mass (EF) of the robot. EF is a property of the robot (each point of the robot), which can be computed using the dynamics of the robot [16]. EF is always computed in a given direction. We consider this direction to be the direction of the link on which the collision is detected. Using this direction , one can compute the EF as
@@
This property of robots is well known in the community, but is arguably still not used often. The reason is its sensitivity to proper dynamic model of the robot and instability in some configurations (mainly near the base of the robot; see [17] for discussion on the topic). We provide some examples in Fig. 4, where the EF in every configuration of our task is shown for three skin pads (the pads that can be in collision with any of the obstacles). We can see that the EF is in most cases lower than HMM, which was the motivation to use this property in our dynamic threshold setting policy. However, we can also see that it is higher in the beginning of the task for one of the pads. The given skin part (grey pad in Fig. 1) is in the beginning of the movement very close to the robot base and is moving in such a way that it realistically cannot exert any force to obstacles (see the video at https://youtu.be/UyOa--6Jpj4).
Setting collision force thresholds. For both of the dynamic collision handling strategies, on collision of a particular robot link, the collision force is estimated based on Eq. 2 and compared to force thresholds obtained from [6]: and for quasi-static and transient contacts, respectively.
The collision reaction is thus defined as:
@@
II-D Transient Contact Force Estimation
In physical HRI, safety is at stake and we need to make sure that the dynamic estimates of the impact forces, which are used to boost productivity here, are accurate or conservative (actual collision forces are always lower than estimated). In our setup, the certified force-measuring device is used only at the location of the quasi-static contact. In order to empirically evaluate also the forces of the transient collisions with the buckets suspended from the ceiling, we measured their deflection after impact and approximated them as pendulum. Therefore, we added a large protractor (circa A4 paper) behind the first bucket B1 and captured the swing of the bucket using a camera and extracted the maximal angle of the pendulum . We know that potential energy and kinetic energy , where is the weight of the bucket, is the length of the string, is the angular velocity and height . Using the law of conservation of energy, the energies at the maximal angle should be equal, i.e., , and from there we can solve for as
For this calculation, we used the baseline, FACTORY, setting where the robot stops immediately after impact. We assume that all momentum will be transferred to the bucket, i.e., , with momentum of the bucket given by . From impulse-momentum principle we know that the average force over a small period of time is equal to the moment, i.e., . If we put all together and solve for we get
III Experiments and Results